# Theory of Valley Splitting in Si/SiGe Spin-Qubits: Interplay of Strain, Resonances and Random Alloy Disorder

### v2

@@ Appendix B Evaluation of Valley Splitting Components and Band Structure Coefficients
In the small strain limit, the reciprocal lattice vectors are written as $\mathbf{G}_{n}=\sum_{j=1}^{3}n_{j}\left(I-\varepsilon\right)\mathbf{b}_{j}$, where the $\mathbf{b}_{j}$ are the primitive reciprocal lattice vectors of the relaxed fcc lattice [56]
@@
Therefore, the difference between two reciprocal lattice vectors reads
@@
We will now evaluate the contribution to the valley splitting from Eq. (24). Evaluation of Eqs. (25) and (34) is carried out along the same lines. The summation in Eq. (24) can be rewritten as
$\Delta_{\mathrm{det}}^{l}$ $=\sum_{\Delta n_{1}}\sum_{\Delta n_{2}}\sum_{\Delta n_{3}}\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\delta_{\mathbf{G}-\mathbf{G}^{\prime},\sum_{j}\Delta n_{j}\mathbf{b}_{j}^{\varepsilon}}$
@@
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-i\left(\sum_{j}\Delta n_{j}\mathbf{e}_{z}^{T}\mathbf{b}_{j}^{\varepsilon}+2k_{0}\right)z}\times$
[…]
where we have inserted a three-dimensional Kronecker delta and summation with $\Delta n_{i=1,2,3}$ running over all integers. As the in-plane extension of the QD is much larger than the lattice constant $l_{x},l_{y}\gg a_{0}$, the Gaussian terms (second line) effectively reduce to simple selection rules encoded by new Kronecker delta functions:
@@
$\mathrm{e}^{-\left(\frac{l_{x}}{2}\sum_{j}\Delta n_{j}\mathbf{e}_{x}^{T}\mathbf{b}_{j}^{\varepsilon}\right)^{2}}$ $\approx\delta_{-\Delta n_{1}+\Delta n_{2}+\Delta n_{3},0}\mathrm{e}^{-\left(\frac{l_{x}}{2}\sum_{j}\Delta n_{j}\mathbf{e}_{x}^{T}\varepsilon\mathbf{b}_{j}\right)^{2}},$
@@
$\mathrm{e}^{-\left(\frac{l_{y}}{2}\sum_{j}\Delta n_{j}\mathbf{e}_{y}^{T}\mathbf{b}_{j}^{\varepsilon}\right)^{2}}$ $\approx\delta_{+\Delta n_{1}-\Delta n_{2}+\Delta n_{3},0}\mathrm{e}^{-\left(\frac{l_{y}}{2}\sum_{j}\Delta n_{j}\mathbf{e}_{y}^{T}\varepsilon\mathbf{b}_{j}\right)^{2}}.$
These selection rules greatly simplify the evaluation of the summation in $\Delta_{\mathrm{det}}^{l}$, which (after renaming the remaining summation index to $n$) yields the compact result
$\Delta_{\mathrm{det}}^{l}=\sum_{n\in\mathbb{Z}}C_{n}^{\left(2\right)}J_{n}^{l}.$ (56)
Here we have introduced the family of integrals indexed by $n\in\mathbb{Z}$
@@
$J_{n}^{l}$ $=\mathrm{e}^{-\frac{1}{4}\left(nG_{0,x}l_{x}\right)^{2}}\mathrm{e}^{-\frac{1}{4}\left(nG_{0,y}l_{y}\right)^{2}}\times$ (57)
@@
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-i\left(nG_{0,z}+2k_{0}\right)z}\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\psi_{0}^{2}\left(z\right)$
and the band structure coefficients
$C_{n}^{\left(2\right)}=\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\delta_{\mathbf{G}-\mathbf{G}^{\prime},n\mathbf{G}_{0}}.$ (58)
Both expressions involve (components of) the displacement vector
@@
$\mathbf{G}_{0}$ $=\mathbf{b}_{1}^{\varepsilon}+\mathbf{b}_{2}^{\varepsilon}=\frac{4\pi}{a_{0}}\left(\begin{array}[]{c}-\varepsilon_{z,x}\\ -\varepsilon_{y,z}\\ 1-\varepsilon_{z,z}\end{array}\right),$ (62)
@@
which is the separation of the two low-energy $X$ points in the strained diamond lattice. The vector $\mathbf{G}_{0}$ determines both the resonance conditions in Eq. (57) as well as the selection rule in Eq. (58). We note that shear strain components $\varepsilon_{z,x}$ and $\varepsilon_{y,z}$ lead to a Gaussian damping in Eq. (57) and are therefore expected to reduce the magnitude of the valley splitting. The compact form of Eq. (56) allows for an efficient numerical evaluation, where the double-summation in Eq. (24) was effectively replaced by a single summation over $J_{n}^{l}$ (with precomputed $C_{n}^{\left(2\right)}$) for a few integers only.
Along the same lines, we obtain for Eq. (25) and (34)
$\Delta_{\mathrm{det}}^{t}$ $=\sum_{n\in\mathbb{Z}}C_{n}^{\left(2\right)}J_{n}^{t}$ (63)
@@
with integrals
$J_{n}^{t}$ $=\mathrm{e}^{-\frac{1}{4}\left(nG_{0,x}l_{x}\right)^{2}}\mathrm{e}^{-\frac{1}{4}\left(nG_{0,y}l_{y}\right)^{2}}\times$ (65)
×$\phantom{=}\times\bigg{(}\frac{\hbar\omega_{x}}{4}\left[1-\left(\frac{nG_{0,x}l_{x}}{\sqrt{2}}\right)^{2}\right]$
$+\frac{\hbar\omega_{y}}{4}\left[1-\left(\frac{nG_{0,y}l_{y}}{\sqrt{2}}\right)^{2}\right]\bigg{)}$
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-i\left(nG_{0,z}+2k_{0}\right)z}\psi_{0}^{2}\left(z\right)$
@@
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-inG_{0,z}z}X\left(z\right)\left(1-X\left(z\right)\right)\psi_{0}^{4}\left(z\right)$
and a second set of band structure coefficients governing the magnitude of the disorder-induced contribution
$C_{n}^{\left(4\right)}$ $=\sum_{\mathbf{G},\mathbf{G}^{\prime}}\sum_{\mathbf{K},\mathbf{K}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)c_{+}\left(\mathbf{K}\right)c_{-}^{*}\left(\mathbf{K}^{\prime}\right)\times$
[…]
@@
Finally, we combine the two deterministic contributions (57) and (65) into a single expression as
$J_{n}^{\mathrm{det}}=J_{n}^{\mathrm{det},t}+J_{n}^{\mathrm{det},l}.$ (68)
Appendix C Statistics
@@
C.1 Intervalley Coupling Parameter
We consider the random component of the intervalley coupling parameter given in Eq. (32) and seek for a characterization of its statistical properties in terms of its characteristic function. The characteristic function is the Fourier transform of the probability density function, which is given for a complex-valued variable as
@@ C.1 Intervalley Coupling Parameter
$w_{j}=\Delta E_{c}\Omega_{a}\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\mathrm{e}^{-i\left(\mathbf{G}-\mathbf{G}^{\prime}+2\mathbf{k}_{0}\right)\cdot\mathbf{R}_{j}}\Psi_{0}^{2}\left(\mathbf{R}_{j}\right).$
@@
$\varphi_{\Delta_{\mathrm{rand}}}\left(s\right)$ $=\prod_{j}\mathrm{e}^{-i\mathrm{Re}\left(s^{*}w_{j}\right)X\left(\mathbf{R}_{j}\right)}\langle\mathrm{e}^{i\mathrm{Re}\left(s^{*}w_{j}\right)N_{j}}\rangle$
@@
$\varphi_{\Delta_{\mathrm{rand}}}\left(s\right)$ $\approx\prod_{j}\left(1-\frac{1}{2}X\left(\mathbf{R}_{j}\right)\left(1-X\left(\mathbf{R}_{j}\right)\right)\left(\mathrm{Re}\left(s^{*}w_{j}\right)\right)^{2}\right)$
$\approx\mathrm{e}^{-\frac{1}{2}\sum_{j}X\left(\mathbf{R}_{j}\right)\left(1-X\left(\mathbf{R}_{j}\right)\right)\left(\mathrm{Re}\left(s^{*}w_{j}\right)\right)^{2}}$
$=\mathrm{e}^{-\frac{1}{4}\left|s\right|^{2}\Gamma-\frac{1}{4}\mathrm{Re}\left(s^{*2}C\right)}.$
This is the characteristic function of a scalar complex-valued normal distribution with covariance $\Gamma$ and relation parameter (or pseudo-covariance) $C$. This is essentially the statement of the central limit theorem for complex-valued random variables. Hence
@@
$\Gamma$ $=\sum_{j}X\left(\mathbf{R}_{j}\right)\left(1-X\left(\mathbf{R}_{j}\right)\right)\left|w_{j}\right|^{2}$ (69)
$=\left(\Delta E_{c}\right)^{2}\Omega_{a}\sum_{\mathbf{G},\mathbf{G}^{\prime}}\sum_{\mathbf{K},\mathbf{K}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)c_{+}\left(\mathbf{K}\right)c_{-}^{*}\left(\mathbf{K}^{\prime}\right)$
×$\phantom{=}\times\int\mathrm{d}^{3}r\,X\left(z\right)\left(1-X\left(z\right)\right)\mathrm{e}^{-i\left(\mathbf{G}-\mathbf{G}^{\prime}-\mathbf{K}+\mathbf{K}^{\prime}\right)\cdot\mathbf{r}}\Psi_{0}^{4}\left(\mathbf{r}\right)$
$=\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle$
@@
$C$ $=\sum_{j}X\left(\mathbf{R}_{j}\right)\left(1-X\left(\mathbf{R}_{j}\right)\right)w_{j}^{2}$ (70)
$=\left(\Delta E_{c}\right)^{2}\Omega_{a}\sum_{\mathbf{G},\mathbf{G}^{\prime}}\sum_{\mathbf{K},\mathbf{K}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)c_{+}^{*}\left(\mathbf{K}\right)c_{-}\left(\mathbf{K}^{\prime}\right)$
×$\phantom{=}\times\int\mathrm{d}^{3}r\,X\left(z\right)\left(1-X\left(z\right)\right)\mathrm{e}^{-i\left(\mathbf{G}-\mathbf{G}^{\prime}+\mathbf{K}-\mathbf{K}^{\prime}+4\mathbf{k}_{0}\right)\cdot\mathbf{r}}\Psi_{0}^{4}\left(\mathbf{r}\right)$
$=\langle\Delta_{\mathrm{rand}}^{2}\rangle.$
@@
we conclude that the real and imaginary parts of $\Delta_{\mathrm{rand}}$ are in fact weakly correlated and have a slightly different variance:
@@
Due to the rapidly oscillating term proportional to $4k_{0}$ in Eq. (70), however, the relation parameter $C$ is typically orders of magnitude smaller than the covariance $\Gamma$, such that $C\approx 0$ is a good approximation. In this limit, the distribution of $\Delta_{\mathrm{rand}}$ becomes circularly symmetric in the complex plane. Including the shift due to the deterministic contribution $\Delta_{\mathrm{det}}=\left|\Delta_{\mathrm{det}}\right|\exp{\left(i\Theta\right)}$, we finally arrive at
@@
$\mathrm{Im}\left(\Delta\right)$ $\sim\mathrm{Normal}\left(\mu=\left|\Delta_{\mathrm{det}}\right|\sin{\left(\Theta\right)},\sigma^{2}=\frac{1}{2}\Gamma\right).$ (72)
C.2 Valley Splitting
The statistical distribution of the valley splitting $E_{\mathrm{VS}}=2\left|\Delta\right|$ is again obtained via computation of its characteristic function. We consider

### v3

@@ Appendix C Evaluation of Valley Splitting Components and Bloch Factor Coefficients
In the small strain limit, the reciprocal lattice vectors are written as $\mathbf{G}_{n}\simeq\sum_{j=1}^{3}n_{j}\left(I-\varepsilon\right)\mathbf{b}_{j}$, where the $\mathbf{b}_{j}$ are the primitive reciprocal lattice vectors of the relaxed fcc lattice [58]
@@
Therefore, the difference between two (strained) reciprocal lattice vectors reads
@@
We will now evaluate the contribution to the valley splitting from Eq. (24). The summation in Eq. (24) can be written as
$\Delta_{\mathrm{det}}$ $=\sum_{\Delta n_{1}}\sum_{\Delta n_{2}}\sum_{\Delta n_{3}}\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\delta_{\mathbf{G}-\mathbf{G}^{\prime},\sum_{j}\Delta n_{j}\mathbf{b}_{j}^{\varepsilon}}$
@@
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-i\left(\sum_{j}\Delta n_{j}\mathbf{e}_{z}^{T}\mathbf{b}_{j}^{\varepsilon}+2k_{0}\right)z}\,\bigg{[}U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)$
@@
$\phantom{=}+\frac{\hbar\omega_{x}}{2}\left(\frac{1}{2}-\left(\frac{l_{x}}{2}\sum_{j}\Delta n_{j}\mathbf{e}_{x}^{T}\mathbf{b}_{j}^{\varepsilon}\right)^{2}\right)$
@@
$\phantom{=}+\frac{\hbar\omega_{y}}{2}\left(\frac{1}{2}-\left(\frac{l_{y}}{2}\sum_{j}\Delta n_{j}\mathbf{e}_{x}^{T}\mathbf{b}_{j}^{\varepsilon}\right)^{2}\right)\bigg{]}\left|\psi_{0}\left(z\right)\right|^{2},$
where we have inserted a three-dimensional Kronecker delta and summation with $\Delta n_{i=1,2,3}$ running over all integers. As the in-plane extension of the QD is much larger than the lattice constant $l_{x},l_{y}\gg a_{0}$, the Gaussian terms (second line) effectively reduce to simple selection rules encoded by Kronecker delta symbols:
@@
$\mathrm{e}^{-\left(\frac{1}{2}l_{x}\sum_{j}\Delta n_{j}\mathbf{e}_{x}^{T}\mathbf{b}_{j}^{\varepsilon}\right)^{2}}$ $\approx\delta_{-\Delta n_{1}+\Delta n_{2}+\Delta n_{3},0}\times$
@@
×$\phantom{\approx}\times\mathrm{e}^{-\left(\frac{1}{2}l_{x}\sum_{j}\Delta n_{j}\mathbf{e}_{x}^{T}\varepsilon\mathbf{b}_{j}\right)^{2}},$
@@
$\mathrm{e}^{-\left(\frac{1}{2}l_{y}\sum_{j}\Delta n_{j}\mathbf{e}_{y}^{T}\mathbf{b}_{j}^{\varepsilon}\right)^{2}}$ $\approx\delta_{+\Delta n_{1}-\Delta n_{2}+\Delta n_{3},0}\times$
×$\phantom{\approx}\times\mathrm{e}^{-\left(\frac{1}{2}l_{y}\sum_{j}\Delta n_{j}\mathbf{e}_{y}^{T}\varepsilon\mathbf{b}_{j}\right)^{2}}.$
These selection rules greatly simplify the evaluation of the summation in $\Delta_{\mathrm{det}}$, which yields the compact result
$\Delta_{\mathrm{det}}=\sum_{n\in\mathbb{Z}}C_{n}^{\left(2\right)}J_{n}^{\mathrm{det}}.$ (48)
Here we have introduced the family of integrals
$J_{n}^{\mathrm{det}}$ $=\mathrm{e}^{-\left(\frac{1}{2}nG_{0,x}l_{x}\right)^{2}}\mathrm{e}^{-\left(\frac{1}{2}nG_{0,y}l_{y}\right)^{2}}$ (49)
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-i\left(nG_{0,z}+2k_{0}\right)z}\,\bigg{[}U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)$
$\phantom{\times=}+\frac{\hbar\omega_{x}}{2}\left(\frac{1}{2}-\left(\frac{nG_{0,x}l_{x}}{2}\right)^{2}\right)$
$\phantom{\times=}+\frac{\hbar\omega_{y}}{2}\left(\frac{1}{2}-\left(\frac{nG_{0,y}l_{y}}{2}\right)^{2}\right)\bigg{]}\left|\psi_{0}\left(z\right)\right|^{2},$
the coefficients
$C_{n}^{\left(2\right)}=\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\delta_{\mathbf{G}-\mathbf{G}^{\prime},n\mathbf{G}_{0}}$ (50)
and the vector
$\mathbf{G}_{0}$ $=\left(I-\varepsilon\right)\left(\mathbf{b}_{1}+\mathbf{b}_{2}\right)=\frac{4\pi}{a_{0}}\left(\begin{array}[]{c}-\varepsilon_{z,x}\\ -\varepsilon_{y,z}\\ 1-\varepsilon_{z,z}\end{array}\right).$ (54)
@@
The vector $\mathbf{G}_{0}$ describes the separation of the two low-energy $X$ points in the strained fcc lattice and determines both the resonance conditions in Eq. (49) as well as the selection rule in Eq. (50). We note that shear strain components $\varepsilon_{z,x}$ and $\varepsilon_{y,z}$ lead to a Gaussian damping in Eq. (49) and are therefore expected to reduce the magnitude of the valley splitting. The compact form of Eq. (48) allows for an efficient numerical evaluation, where the double-summation in Eq. (24) was effectively replaced by a single summation over $J_{n}^{\mathrm{det}}$ (with precomputed $C_{n}^{\left(2\right)}$) for a few integers only.
Along the same lines, we obtain for Eq. (33)
@@
with the integral
@@
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-inG_{0,z}z}X\left(z\right)\left(1-X\left(z\right)\right)\left|\psi_{0}\left(z\right)\right|^{4}$
and a second set of coefficients governing the magnitude of the disorder-induced contribution
$C_{n}^{\left(4\right)}=\sum_{m}C_{m-n}^{\left(2\right)*}C_{m}^{\left(2\right)}.$ (57)
@@
Appendix D Statistics
@@
D.1 Intervalley Coupling Parameter
We consider the random component of the intervalley coupling parameter given in Eq. (31) and seek a characterization of its statistical properties in terms of its characteristic function. The characteristic function is the Fourier transform of the probability density function, which is given for a complex-valued variable as
@@ D.1 Intervalley Coupling Parameter
$w_{j}$ $=\Delta E_{c}\Omega_{a}\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\times$
[…]
@@
$\varphi_{\Delta_{\mathrm{rand}}}\left(s\right)$ $=\prod_{j}\mathrm{e}^{-i\mathrm{Re}\left(s^{*}w_{j}\right)X\left(\mathbf{R}_{j}\right)}\langle\mathrm{e}^{i\mathrm{Re}\left(s^{*}w_{j}\right)N_{j}}\rangle,$
@@
$\varphi_{\Delta_{\mathrm{rand}}}\left(s\right)\approx$
$\quad\approx\prod_{j}\left(1-\frac{1}{2}X\left(\mathbf{R}_{j}\right)\left(1-X\left(\mathbf{R}_{j}\right)\right)\left(\mathrm{Re}\left(s^{*}w_{j}\right)\right)^{2}\right)$
$\quad\approx\mathrm{e}^{-\frac{1}{2}\sum_{j}X\left(\mathbf{R}_{j}\right)\left(1-X\left(\mathbf{R}_{j}\right)\right)\left(\mathrm{Re}\left(s^{*}w_{j}\right)\right)^{2}}$
$\quad=\mathrm{e}^{-\frac{1}{4}\left|s\right|^{2}\Gamma-\frac{1}{4}\mathrm{Re}\left(s^{*2}C\right)}.$
This is the characteristic function of a scalar complex-valued normal distribution with covariance $\Gamma$ and pseudo-covariance $C$. Hence
@@
$\Gamma=\sum_{j}X\left(\mathbf{R}_{j}\right)\left(1-X\left(\mathbf{R}_{j}\right)\right)\left|w_{j}\right|^{2}=\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle$ (58)
@@
$C=\sum_{j}X\left(\mathbf{R}_{j}\right)\left(1-X\left(\mathbf{R}_{j}\right)\right)w_{j}^{2}=\langle\Delta_{\mathrm{rand}}^{2}\rangle.$ (59)
@@
we conclude that the real and imaginary parts of $\Delta_{\mathrm{rand}}$ are weakly correlated and have a slightly different variance:
@@
Due to the rapidly oscillating term proportional to $4k_{0}$ in Eq. (59), however, the relation parameter $C$ is typically orders of magnitude smaller than the covariance $\Gamma$, such that $C\approx 0$ is a good approximation. In this limit, the distribution of $\Delta_{\mathrm{rand}}$ becomes circularly symmetric in the complex plane. Including the deterministic contribution $\Delta_{\mathrm{det}}=\left|\Delta_{\mathrm{det}}\right|\exp{\left(i\Theta\right)}$, we finally arrive at
@@
$\mathrm{Im}\left(\Delta\right)$ $\sim\mathrm{Normal}\left(\mu=\left|\Delta_{\mathrm{det}}\right|\sin{\left(\Theta\right)},\sigma^{2}=\frac{1}{2}\Gamma\right),$ (61)
where $\Theta$ is the deterministic valley phase.
D.2 Valley Splitting
The statistical distribution of the valley splitting $E_{\mathrm{VS}}$ is again obtained via computation of its characteristic function. We consider